OpenStudy (anonymous):

why doesnt x^3-8=(x-2)^3??? please reply!

OpenStudy (amistre64):

becasue they are 2 different operations

OpenStudy (amistre64):

exponents are(nt) distributive across addition

OpenStudy (anonymous):

In one you are cubing the x and subtracting the 8. In the other you are subtracting the 2 then cubing it. On a side not: I WANNA BE A MOD AMISTRE!!!

OpenStudy (anonymous):

because (x-2)^3 is like (x-2)(x-2)(x-2)

OpenStudy (anonymous):

so, it doesn't equal the same as x^3-8

OpenStudy (anonymous):

AMISTRE!!! D: My lip is poking out :(

OpenStudy (anonymous):

Multiply out \[(x-2)^3\] and you will see that the answer is not \[x^3-8\]

OpenStudy (anonymous):

In general: \[(a+b)^n \neq a^n+b^n\]

OpenStudy (anonymous):

but how would u know it is (x-2)(x^2+2x+4)

OpenStudy (anonymous):

We agree that 2^3=2*2*2? So: (x-2)^3=(x-2)(x-2)(x-2) So you can get: (x-2)(x^2+2x+4)

OpenStudy (anonymous):

ohhh i understand thanks so muh guys!